INAF queue observing, 2015 Mar 20-21

Observer: R. Speziali, S. Bisogni
Telescope Operator: D. Gonzales Huerta
Support Astronomer: Olga on call
Instrument: MODS - LBC-R

Summary

At the sunset the sky is not so covered, but the mountain is wrapped in a big cloud passing. Humidity is 100.3%. The dome stays closed. Some chances for later in the night.
UT 03:05 TO opens the dome. He is trying to guiding. Seeing 2.3''.

Night duration (12\xB0 el twilight) 19:30-05:30 = 11 hrs

Observing time: 2.5 hrs

Weather time loss: 8.5 hrs (~77%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Saracco C RDCS red grating 1h seeing 2.5'' out of spec
Fiore A goodsnorth1 dual grating + i-sloan filter (LBC-R) 1.5h average seeing 1.2''
MODS Calibrations (bias)        
02:30 Stay Close

03:05 Opening

03:35 start acquisition of Saracco 19C

03:35 problem with the telescope

03:45 start acquisition again. Seeing=2" (out of spec)

03:55 execMODS mods.11.RDCSm2_2.obs (seeing=2.5") mods1r.20150321.0004-6

04:36 Acquisition of FIORE B

04:37 telescope error. Recollimating

04:45 Start Acquisition again (seeing =2" - out of spec) mods1r.20150321.0007-0009

04:46 DOFPIA

04:55 We are close to close the dome due to high Humidity

05:11 execMODS goodsnorth2_mos2.obs (three exposures left)
  • we are trying to collimate with LBC-R but DOFPIA doesn't converge. The TO calls the support astronomer who is analyzing the problem.
  • seeing 1.6'' during the first exposure

05:19 The TO closes the dome, we have to abort the MODS integration.

06:05 the TO finds some problems opening the dome; one door is stuck

06:25 the TO opens the dome, slewing to Fiore A

06:28 acqMODS goodsnorth1_mos.acq mods1.20150321.0010-0012
  • DOFPIA
  • average seeing 2.3''

06:46 acqMODS goodnorth1_mos.acq; we acquire again, in one of the boxes (the 3rd) the alignment star is missing. mods1r.20150321.0013-0015; We contact Olga. She could find it even if it was very faint.

07:00 execMODS goodsnorth1_mos_1.obs mods1r.20150321.0016-0019 mods1b20150321.0002-0005
  • seeing 1.3'' during the first MODS integration
  • LBC images seem good
  • seeing 1.2" - WE'R RIDING HIGH ... FINALLY!
  • seeing 1.2'' during the 2nd integration
  • 07:27 we run DOFPIA again
  • seeing 1.1'' at the beginning of the 3rd integration
  • seeing 1.2'' at the beginning of the 4th integration

08:12 execMODS goodsnorth1_mos_2.obs mods1r.20150321.0020-0021 mods1b.20150321.0006-0007
  • seeing 1.1'' at the beginning of the first integration
  • seeing 1.3'' at the beginning of the 2nd integration
  • 08:41 (2nd integration) the seeing is getting worse (1.4'')

08:47 the TO closes the dome once we have completed the 2nd integration

We run ./doBias mods1r.20150321.0022-0031 mods1b.20150321.0008-0017

Comments


 

-- %USERSIG{SusannaBisogni - 2015-03-21}%
Topic revision: r13 - 29 Apr 2015, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback